OpenStudy (bekkah323):

Paint town sold 45 paintbrushes, one kind at $8.50 each and another at $9.75 each. in all, $398.75 was taken in for the brushes. how many of each kind were sold.

OpenStudy (anonymous):

It's a system of equations

OpenStudy (nurali):

Let x be number of $8.50 brushes; y be number of $9.75 brushes then x+y = 45------(1) 8.5x+9.75y= 398.75----(2) Multiplying first equation by 8.5 and subtracting gives 1.25y=16.25 y=16.25/1.25 y=13 put 'y' in equation (1) x=32 So there were 32 paintbrushes costing $8.50 and 13 paintbrushes costing $9.75